PSEUDO-r_{e} STRUCTURES FROM EXPERIMENTAL ROTATIONAL CONSTANTS AND AB INIT IO VIBRATION-ROTATION CONSTANTS

\underline{P. GRONER}, and R. D. WARREN, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110.

The determination of accurate molecular structures from spectroscopic data has been and remains a formidable challenge. Inspired by the close agreement of experimental and theoretical centrifugal distortion constants and inertial defects ^{a}, we explored the possibility of using vibration-rotation constants from ab initio calculations together with experimental ground state constants to derive approximate r_{e} structures. In the past, this has been done successfully for (very) small molecules with large basis sets at high levels of calculation (for an example, see ^{b}). We were interested in finding out whether MP2 calculations with small basis sets yield sufficiently accurate vibration-rotation constants for larger molecules. This method was tested with MP2/6-31G(d) calculations for the COCl_{2}, HCOCl, FCOCl, and HCOOH, for which near equilibrium structures derived from experimental data are available in the literature. It was subsequently used to reevaluate the structures of CF_{2}NH, CF_{e}NCl, CH_{3}OCOCl, CH_{3}OCOCN, and two conformers of CH_{3}CH_{2}PH_{2}. Many, but not all, problems encountered in deriving meaningful structures are significantly reduced. It appears that the structures are comparable to r_{m}^{ρ} ^{c} and r_{m}^{(1)}/r_{m}^{(2)} ^{d} structures.

aP. Groner, J. R. Durig, D. D. DesMarteau, and S.-H. Hwang, J. Chem. Phys. 110, 9411 (1999). P. Groner, J. R. Durig, and D. D. DesMarteau, J. Chem. Phys. 105, 7263 (1996).

bM. Oswald, J. Flügge, and P. Botschwina, J. Mol. Struct. 320, 227 (1994).

cM. D. Harmony, in: J. R. Durig (Ed.), Vibrational Spectra and Structure, (Elsevier, Amsterdam, 2000) Vol. 24.

dJ. K. G. Watson, A. Roytburg, and W. Ulrich, J. Mol. Spectrosc. 196, 102 (199).